John traveled north from Miami at an average speed of 50mph. Two hours after he left, Sharon left from Miami traveling in the same direction at an average speed of 55 mph. HOw long will it take Sharon to catch up with John?
Answer by richwmiller(17219)

Gaining speed 55-50= 5 mph
Head start 50*2 =100 miles
Time to catch up 100/5=20 hours
Algebraic solution
50(t+2)=55t
50t+100=55t
100=55t-50t
100=5t
t=20 hours
Distance to meeting 20*55=1100 miles
code drtch1